### v2

# Analysis of reactor burnup simulation uncertainties for antineutrino spectrum prediction

Abstract
Nuclear reactors are a source of electron antineutrinos due to the presence of unstable fission products that undergo $\beta^{-}$ decay. They will be exploited by the JUNO experiment to determine the neutrino mass ordering and to get very precise measurements of the neutrino oscillation parameters. This requires the reactor antineutrino spectrum to be characterized as precisely as possible both through high resolution measurements, as foreseen by the TAO experiment, and detailed simulation models. In this paper we present a benchmark analysis utilizing Serpent Monte Carlo simulations in comparison with real pressurized water reactor spent fuel data. Our objective is to study the accuracy of fission fraction predictions as a function of different reactor simulation approximations. Then, using the BetaShape software, we construct fissile antineutrino spectra using the summation method, thereby assessing the influence of simulation uncertainties on the reactor antineutrino spectrum.
Keywords:
Reactor antineutrino spectra Fission fractions Fuel burnup Summation method
journal: -
1 Introduction
The knowledge of the electron antineutrino ($\bar{\nu}_{e}$) flux emitted by nuclear reactors is of crucial importance for the Jiangmen Underground Neutrino Observatory (JUNO) An2016 ; JUNO_new_ref experiment, which will exploit 8 reactors as $\bar{\nu}_{e}$ sources. The JUNO detector, a 20-thousand-ton liquid scintillator located at $\sim$53 km from the Yangjiang and Taishan Nuclear Power Plants, is in its final construction phase in China.
The reactor $\bar{\nu}_{e}$ will be detected through the Inverse Beta Decay (IBD) reaction:
$\bar{\nu}_{e}+p\xrightarrow[]{}n+e^{+}$ (1)
which has a $\bar{\nu}_{e}$ energy threshold at 1.8 MeV.
One of the main goals of the JUNO experiment is to determine the neutrino mass ordering by analyzing the pattern of $\bar{\nu}_{e}$ oscillations at medium baseline. This experimental measurement requires high accuracy, therefore it is mandatory to control all the systematic uncertainties related to the reactor $\bar{\nu}_{e}$ source. For this purpose, the Taishan Antineutrino Observatory (TAO) TAO detector will measure with unprecedented energy resolution the non-oscillated $\bar{\nu}_{e}$ spectrum at $\sim 30$ m from one of the reactors exploited by JUNO. However, since the reactor $\bar{\nu}_{e}$ spectrum changes as a function of reactor fuel burnup, computational tools are needed to assess the $\bar{\nu}_{e}$ flux emitted by all reactors.
Until now, two different methods have been used to model the reactor $\bar{\nu}_{e}$ flux: the conversion Huber2011 ; Mueller2011 and the summation (also called ab-initioFallot2012 ; Estienne_ab_initio methods. Both approaches are affected by uncertainties related to the input data, leading to a discrepancy in both $\bar{\nu}_{e}$ flux intensity and spectral shape HayesVogel2016 ; zhang2023reactor when compared with the experimental data acquired by near detectors DayaBay:2022jnn ; DoubleChooz:2019qbj ; RENO:2015ksa ; Stereo2023 ; Ko_2017 ; Andriamirado_2023 .
Since JUNO and TAO will feature an unprecedented energy resolution in the field of $\bar{\nu}_{e}$ detectors, the summation method is the most suitable to use, because it gives access to the fine structure of the spectrum Dwyer2015 .
To determine the real time reactor $\bar{\nu}_{e}$ flux and spectrum emitted by a nuclear reactor two ingredients, completely independent of each other, are needed:
  1. 1.
    the concentration of $\beta^{-}$ decaying nuclides in the reactor core as a function of time;
  2. 2.
    the $\bar{\nu}_{e}$ spectrum characteristics of each $\beta^{-}$ decaying nuclide.
In this paper we present the development of an analysis tool designed explicitly for probing the uncertainties inherent in the first ingredient. We choose as test model a standard PWR system due to the data availability of most of its components. Even if the reactor design is not strictly equal to the reactors placed in Taishan and Yangjiang, the presented results are representative of a light water reactor burnup cycle. This work is meant to show the methodological approach to reproduce the antineutrino spectrum from a reactor model, highlighting some features that can have an impact on it. Future works will be devoted to have an high fidelity simulation of both the plants involved in JUNO and TAO. Through meticulous simulation of reactor neutronics and fuel burnup, we attain a comprehensive understanding of the production and decay processes of all contributing nuclides that shape the $\bar{\nu}_{e}$ flux. In a second step we couple reactor burnup simulations to an extensive $\bar{\nu}_{e}$ spectral database to investigate the reactor-related uncertainties affecting the non-oscillated reactor $\bar{\nu}_{e}$ spectra. This aspect can be considered of extreme importance for all reactor neutrino experiments since it gives the key to reactor $\bar{\nu}_{e}$ flux control, allowing its real time spectral reconstruction and providing a continuous monitor of the $\bar{\nu}_{e}$ flux variations.
This paper is organised as follows:
  • in Sect. 2 we describe the methodology and the tools that we use to perform our analysis;
  • in Sect. 3 we perform a benchmark analysis with the experimental data of a power reactor, to validate the Monte Carlo simulations of fuel burnup at different levels of approximation;
  • in Sect. 4 we compare the fission fractions and the non oscillated reactor $\bar{\nu}_{e}$ spectrum at different burnup levels obtained with different approximations in the reactor simulations.
2 Methodology and tools
[…]
The resulting spectrum of $\bar{\nu}_{e}$ emitted per unit time by a nuclear reactor $S(E_{\nu},t)$ can be quantified through the following formula:
$S(E_{\nu},t)=\frac{W_{\text{th}}(t)}{\sum_{i=1}^{4}E_{i}f_{i}(t)}\sum_{i=1}^{4}f_{i}(t)S_{i}(E_{\nu},t)$ (2)
where $i$ labels the main fissile nuclides in PWR reactors (${}^{235}$, ${}^{238}$, ${}^{239}$u, ${}^{241}$u), $W_{th}(t)$ is the reactor thermal power, $E_{i}$ is the average energy deposited in the reactor per fission of the $i$-th nuclide, $f_{i}(t)$ are the fission fractions (i.e. the ratio of the fission rate of the $i$-th nuclide to the total fission rate). Finally, $S_{i}(E_{\nu},t)$ is the spectrum associated to the fission of the $i$-th nuclide estimated with the summation method, where the antineutrino spectra of the daughters of $i$ are summed together, each with a weight that reflects their decay rate in the system. It is worth to highlight that the time dependence in $S_{i}(E_{\nu},t)$ will not be considered in this work because of the equilibrium hypothesis (see Sec. 2.2)
Equation (2) highlights that the reactor $\bar{\nu}_{e}$ spectrum depends on two sets of parameters: the first one, including $f_{i}(t)$ and $W_{\text{th}}(t)$, is related to reactor history, power and fuel burnup. The second comprises the nuclear data used to extract the different $\bar{\nu}_{e}$ spectra from the corresponding $\beta^{-}$ decays. In the following subsections, we describe the tools that we use to
  • calculate reactor-related parameters (2.1) using the flexible Serpent environment for fissile system simulations;
  • build the $\bar{\nu}_{e}$ spectra of the four actinides (2.2) adopting the summation method, with the recently developed BetaShape code.
These tools will be subsequently coupled according to the scheme reported in Fig. 1, to get the reactor $\bar{\nu}_{e}$ spectrum.
Reactor spectrum S(E_ν,t) =Wthi=14Eifi(t) ∑_i=1^4 f_i (t) S_i(E_ν) Summation method S_i(E_ν) = ∑_j=1^N_i y_i,j   S_j(E_ν) Fission yields y_i,j Daughters $\bf{\bar{\nu}_{e}}$ spectrumS_j(E) BetaShape Nuclear Database Fission Fractions f_i(t) MC simulation: Serpent Input: Geometry/materials and power history $W_{th}$ Antineutrino dataReactor data
Figure 1: Scheme for the calculation of the reactor antineutrino spectrum. Green box: starting from accessible plant information, the Monte Carlo tool predicts the evolution of fission fractions. The importance of this methodology lies in the flexibility with which input parameters can be varied to accommodate daily power profiles or different reactor designs. Red box: we exploit spectra computed with the BetaShape tool along with the cumulative fission yields to evaluate the fissile spectra at equilibrium with the summation method.
2.1 The Serpent tool for reactor simulations
Neutron-matter interactions occurring within a fissile system (i.e. a nuclear reactor) can be described by the coupling of two equations:
  • Steady State Boltzmann equation BellGlasstone : by knowing the concentration of all nuclides in the system (nuclide field) $N(\mathbf{r},t)$, this equation solves the criticality problem in the state space $\mathbf{s}=(\mathbf{r},\mathbf{\Omega},E,t)$ for the angular neutron flux $\varphi$ with a balance between the neutrons produced by fission and the neutrons which disappear by leakage or absorption:
    $\text{L}(N)\varphi(\mathbf{s})-\frac{1}{k_{\text{eff}}}\text{F}(N)\varphi(\mathbf{s})=0,$ (3)
    where $\text{L}(N)\varphi(\mathbf{s})$ represents the migration and loss of neutrons from $\mathbf{s}$, $\text{F}(N)\varphi(\mathbf{s})$ accounts for the neutron production in $\mathbf{s}$ due to fission and $k_{\text{eff}}$ is the multiplication factor.
  • Bateman equation Bateman : it describes the time evolution of a nuclide field subjected to a neutron flux. This equation (also called burnup equation) takes into account both the fuel depletion and the transmutation processes in the system. The Bateman equation reads as an Ordinary Differential Equation (ODE) system:
    $\frac{\text{d}N(\mathbf{r},t)}{\text{d}t}=\left[\int_{0}^{\infty}\phi(\mathbf{r},E,t)\mathbb{X}(E,T)\,\text{d}E+\mathbb{D}\right]N(\mathbf{r},t)$ (4)
    where $N(\mathbf{r},t)$ is the array of nuclide concentrations, $\phi(\mathbf{r},E,t)=\int_{4\pi}\varphi(\mathbf{s})\,\text{d}\mathbf{\Omega}$ is the scalar neutron flux at time $t$, $\mathbb{X}(E,T)$ is the cross-section (dependent on the neutron energy $E$ and the local material temperature $T$) and fission yield matrix, and $\mathbb{D}$ is the decay matrix, which accounts for the free radioactive evolution.
Many codes are able to solve the neutron transport problem described by Eq. (3) using the Monte Carlo (MC) method. Some MC codes, specialized for reactor simulations, are supplemented by burnup sequence schemes that allow to solve the coupled system described by Eq. (3) and (4). The latter equation has a formal solution only when the neutron flux is constant in time. However, in reality, the neutron flux is a time-dependent quantity because it varies depending on the nuclide field distribution. In order for the analytical solution to be used, the time frame is divided in intervals in which the flux is treated as a time-constant variable. Several methods are given in the literature for choosing the constant value to be used within each interval BurnupMethods .
The three-dimensional continuous-energy Monte Carlo code Serpent SERPENT was chosen for this work. This analysis tool, created by the Finnish research Centre VTT, is capable of simulating the neutronics and fuel evolution of nuclear reactors with a great level of detail, therefore allowing to estimate the neutron flux and, subsequently, derive the reactions rates. In particular, it is possible to compute:
$R=\frac{1}{V}\int_{V}\int_{4\pi}\int_{E_{g-1}}^{E_{g}}h(\mathbf{r},E)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}\mathbf{\Omega}\,\text{d}E$ (5)
where $V$ and $E_{g-1},E_{g}$ are, respectively, the spatial and energy domains in which the neutron flux is estimated, while $h(\mathbf{r},E)$ is a generic response function. The choice of $h$ allows different quantities to be estimated during the neutron transport simulation. In the framework of this work, the most used response functions are:
  • $h=1$ to estimate the volume averaged neutron flux;
  • $h=n_{i}(\mathbf{r})\sigma_{f,i}(E,T)$, where $n_{i}(\mathbf{r})$ is the atomic density of the $i$-th nuclide and $\sigma_{f,i}(E,T)$ its microscopic fission cross section, to estimate the fission reaction rates per unit volume;
  • $h=\sum_{k=1}^{4}n_{k}(\mathbf{r})\sigma_{f,k}(E,T)$ to get the total fission reaction rate per unit volume.
Within the reactor $\bar{\nu}_{e}$ analysis framework, the integration ranges to be used in Eq. (5) to get the fission fractions $f_{i}$ are the whole energy spectrum and the entire spatial domain:
$f_{i}=\frac{\int_{V}\int_{0}^{\infty}n_{i}(\mathbf{r})\sigma_{f,i}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E}{\sum_{k=1}^{4}\int_{V}\int_{0}^{\infty}n_{k}(\mathbf{r})\sigma_{f,k}(E,T)\phi(\mathbf{r},E)\,\text{d}\mathbf{r}\,\text{d}E}$ (6)
2.2 Antineutrino spectra production
As previously pointed out, each fissile nuclide $i$ can be associated with a corresponding $\bar{\nu}_{e}$ spectrum $S_{i}(E_{\nu},t)$ given by the decays of its fission products. The spectrum, normalized, can be written as:
$S_{i}(E_{\nu},t)=\frac{\sum_{j=1}^{N_{i}}A_{i,j}(t)\,S_{j}(E_{\nu})}{\sum_{j=1}^{N_{i}}A_{i,j}(t)\int_{0}^{\infty}\,S_{j}(E_{\nu})\text{d}E_{\nu}}$ (7)
where $A_{i,j}(t)$ is the activity of the nuclide $j$ resulting from the fission of the nuclide $i$, and $S_{j}(E_{\nu})$ is the $\bar{\nu}_{e}$ spectrum associated to the decay of the fission product $j$.
Even if it is possible to get $A_{i,j}(t)$ from Serpent simulations, in this work we decided to build the $\bar{\nu}_{e}$ spectra at equilibrium (i.e. under the condition that the production rate of a nuclide equals its disappearance), thus removing the time dependence in Eq. (7). Since transmutation processes due to neutron capture are relevant only for radionuclides with $\beta^{-}$ Q-value below the IBD energy threshold (1.8 MeV), we can neglect their contribution assuming that, at equilibrium:
$A_{i,j}\propto y_{i,j}$ (8)
where $y_{i,j}$ is the cumulative fission yield (CFY), i.e. the probability of production of nuclide $j$ per each fission of $i$. Moreover, it is worth noting that most part of $\beta^{-}$ decays with Q-values higher than 1.8 MeV are characterized by half-lives much shorter than the typical reactor steady-state operation time, thus the equilibrium condition is quickly reached for most nuclides emitting $\bar{\nu}_{e}$ detectable through the IBD reaction. The adoption of the equilibrium approximation greatly simplifies the computation of $S_{i}(E_{\nu})$, since the values of $y_{i,j}$ and their uncertainties are tabulated in nuclear databases for fission induced by thermal, fast and 14 MeV neutrons.
The other ingredients needed to build the $\bar{\nu}_{e}$ spectrum associated to the $i$-th nuclide are the $\bar{\nu}_{e}$ spectra emitted by its fission products. For this purpose, we decided to exploit the BetaShape code, developed at the Laboratoire National Henri Bequerel (LNHB). This code has been developed to improve the accuracy in the analytical calculations of beta spectra. BetaShape takes as input the Evaluated Nuclear Structure Data Files (ENSDF) published in various nuclear databases and provides as outputs the electron and $\bar{\nu}_{e}$ spectra of any $\beta^{-}$ transitions. More details on the assumptions and approximations adopted in BetaShape can be found in BetaShape_ANewCode .
3 Validation of reactor simulations
In order to validate the Monte Carlo simulations that we use to compute the fission fractions, we performed a benchmark analysis of simulation results with available experimental measurements. Since direct fission fractions measurements are not possible from an operational point of view, we adopt an indirect approach, validating our burnup simulations with published measurements of nuclide concentrations in the spent fuel of a PWR reactor. Indeed, as shown in Eq. 6, the fission fraction $f_{i}(t)$ depends on the concentrations $n_{k}(\mathbf{r})$ of the four fissile nuclides. If we neglect the uncertainties related to the microscopic cross section (which will be addressed in future analyses), it is possible to conclude that the accuracy in the estimation of the fission fractions can be traced back to the ability to reproduce the nuclide field (or the neutron flux, since these two parameters are connected by the burnup scheme).
During the past decades, experimental campaigns have been carried out in order to obtain nuclide concentration data of irradiated fuels, through destructive radiochemical techniques PIE_ORNL . These measurements are considered internationally established and a recognised benchmark. In this work, we use the Takahama-3 data-set as a test case, comparing the concentrations of the four fissile nuclides sampled in two spent fuel rods with the predictions of our Serpent MC simulations. Among the various references in the literature, the Takahama-3 reactor has features very similar, in terms of neutron energy spectrum and fuel type, to those of the reactors that will be used by JUNO and TAO. Moreover, its fuel assemblies reached a high level of burnup, thereby allowing for having a larger isotopic inventory, useful for a time-resilient validation of the simulation outcomes.
Figure 2: Top view of Takahama-3 fuel assembly design. The yellow circle indicates the position of SF95, the pink one shows the SF96 position, whereas the light green is representative of the SF97 location. The red color is used for the other standard fuel pins, whereas dark green is used for the Gd-loaded ones. Finally, blue represents water.
[…]
In Fig. 2 we show the position of the rods SF95 and SF97 in their assemblies and in Table 1 we summarize the available information about the position and the local burnup (estimated through the ${}^{148}$d method Nd148_method1 ; Nd148_method2 ) of the samples taken from the rods labelled with a progressive number from highest to lowest position.
Assembly Sample ID Axial position (cm) Burnup (MWd/kg)
$\bf{{}^{235}}$ (g/tonU)
$\bf{{}^{238}}$ (g/tonU)
$\bf{{}^{239}}$u (g/tonU)
$\bf{{}^{241}}$u (g/tonU)
NT3G23
SF95-1
4.12
14.30
2.674×$\times 10^{4}$
9.499×$\times 10^{5}$
4.227×$\times 10^{3}$
3.69×$\times 10^{2}$
SF95-2
20.2
24.35
1.927×$\times 10^{4}$
9.424×$\times 10^{5}$
5.655×$\times 10^{3}$
9.578×$\times 10^{2}$
SF95-3
72.2
35.53
1.326×$\times 10^{4}$
9.338×$\times 10^{5}$
6.194×$\times 10^{3}$
1.486×$\times 10^{3}$
SF95-4
200.2
36.69
1.23×$\times 10^{4}$
9.335×$\times 10^{5}$
6.005×$\times 10^{3}$
1.466×$\times 10^{3}$
SF95-5
340.2
30.40
1.544×$\times 10^{4}$
9.388×$\times 10^{5}$
5.635×$\times 10^{3}$
1.153×$\times 10^{3}$
NT3G24
SF97-1
0.40
17.69
2.347×$\times 10^{4}$
9.493×$\times 10^{5}$
3.844×$\times 10^{3}$
4.237×$\times 10^{2}$
SF97-2
19.1
30.73
1.571×$\times 10^{4}$
9.377×$\times 10^{5}$
5.928×$\times 10^{3}$
1.235×$\times 10^{3}$
SF97-3
46.8
42.16
1.030×$\times 10^{4}$
9.282×$\times 10^{5}$
6.217×$\times 10^{3}$
1.689×$\times 10^{3}$
SF97-4
168
47.03
8.179×$\times 10^{3}$
9.246×$\times 10^{5}$
6.037×$\times 10^{3}$
1.77×$\times 10^{3}$
SF97-5
276.7
47.25
7.932×$\times 10^{3}$
9.247×$\times 10^{5}$
5.976×$\times 10^{3}$
1.754×$\times 10^{3}$
SF97-6
339.7
40.79
1.016×$\times 10^{4}$
9.310×$\times 10^{5}$
5.677×$\times 10^{3}$
1.494×$\times 10^{3}$
Table 1: Data of axial positions and local burnup of the samples drilled from the SF95 and SF97 pins. The axial locations have to be intended as the distance from the top of the fuel, as reported in the original documentation. The local burnup was evaluated with the ${}^{148}\text{Nd}$ technique, which brings a ±$\pm 3\%$ uncertainty. Finally we report the measured concentrations for ${}^{235}$, ${}^{238}$, ${}^{239}$u and ${}^{241}$u in units of grams per uranium tons in the fresh fuel. Data from Takahama3 .
3.1 Geometry and materials implementation
In order to simulate the fuel evolution in the SF95 and SF97 pins of the Takahama reactor, we implemented the geometry and the materials of the whole assembly in Serpent, taking as reference the first table in Takahama3 . Fig. 2 shows the assembly section in the xy plane of the simulated geometry. At the boundaries of the x and y directions, we apply reflective conditions to simulate the presence of the surrounding fuel assemblies, whereas in the z direction the geometry extends for 4.46 m, including the caps of the fuel cladding and a 20 cm thick water layer at top and bottom, to take into account the reflection of neutrons by water. To define the materials we used the cross sections of the ENDF/B-VII.1 ENDF_B_VII nuclear library.
3.2 Neutronic simulation of fresh fuel
Figure 3: Axial neutron flux of SF95 rod at the beginning of the depletion cycle.
Figure 4: Distribution of neutrons in the energy range in the SF95 rod location at the beginning of the depletion cycle.
The assembly was simulated at the beginning of the depletion cycle in the fresh fuel configuration, with the purpose of studying the spatial and energy distributions of neutron fluxes. In Fig. 3, we show the axial profile of the neutron flux intensity, obtained by setting an infinite energy domain in Eq. (5) and a spatial mesh with 100 equal subdivisions from $z=-182.4$ cm to $z=182.4$ cm, corresponding to the volume occupied by the fuel inside the pin. The bell shape of the neutron flux axial profile has the effect of burning the ${}^{235}$ atoms faster in the central part of the rod. As a consequence, as the fuel depletion proceeds, the axial profile becomes more flat, progressively equalizing the nuclide concentrations along the z axis. In Fig. 4, we show the neutron energy distribution, obtained by setting the spatial domain in Eq. (5) equal to the entire fuel volume and by dividing the energy range into 250 bins from $E=0.001$ eV to $E=20$ MeV (with the same width in logarithmic scale [^1] This subdivision of the energy bins is also called lethargic scale.).
3.3 Power normalization
Start Stop Cycle State Time interval (d) Rods
26/01/1990 15/02/1991 5 On 385 SF95, SF96, SF97
15/02/1991 14/05/1991 Cool 88 SF95, SF96, SF97
14/05/1991 19/06/1992 6 On 402 SF95, SF96, SF97
19/06/1992 20/08/1992 Cool 62 SF97
20/08/1992 30/09/1993 7 On 406 SF97
Table 2: Irradiation history of Takahama-3 experiment. Data taken from Takahama3 .
SF95 SF97
Cycle 5 0.04576 0.037147
Cycle 6 0.04013 0.039128
Cycle 7 - 0.034402
Table 3: Average power densities used to normalize the Serpent simulations of the Takahama-3 fuel burnup. Entries are in kW/gU of initially loaded uranium.
In this work we simulate with a good level of approximation the irradiation history of the two Takahama assemblies, considering both the different power levels during the reactor operation cycles [^2] Each cycle usually has a duration of $\sim 12-18$ months, after which each assembly can either be removed from the core or moved into another core location. and the cooling periods. The technical report cited above (Takahama3 ) provides the power density values at the sampling points as a function of time since reactor start-up (approximately, monthly data). The unit used to express power density is kW/g(U), i.e. kW divided by the mass of Uranium initially loaded in the sampled volume. From these reactor operation data, it is possible to estimate the average power density for every pin in each depletion cycle, so as to set the correct input power values (used to normalize the neutron fluxes and the reaction rates) in the Serpent simulations. In particular, the following steps were performed:
  • an effective local power value was calculated for each sampling point as the average value of the powers in the pin over the time intervals when the reactor was at full power;
  • after extracting the time-weighted axial power profile for the SF95 and SF97 pins, the integral mean value of the power was calculated to get the effective power density for the normalization.
In Table 2 we report the dates and the time intervals in which the assemblies were in operation or in a cooling phase. Fuel assembly NT3G23 (and thus the fuel pin SF95) was irradiated for two cycles, whereas the assembly NT3G24 (and thus the fuel pin SF97) was irradiated for three cycles, reaching a higher burnup level than the previous one. Finally, Table 3 lists the average specific power used for the pins in the cycles.
3.4 Burnup simulation
As introduced in Section 2.1, a burnup simulation implies the coupling of two equations. The first one is the Boltzmann equation, which determines the neutron flux by knowing the material composition of the system (i.e., the isotopic concentrations) and is solved using the Monte Carlo routine. The second one is the Bateman equation, solved through a direct numerical method, which provides the time evolution of the materials in the system when the flux distribution at an initial time is known. These two equations are solved alternately during the depletion chain accordingly to a burnup scheme. The time discretization choice (i.e., how many times to solve the Boltzmann equation) is nontrivial. In this work we divided the time frame into two parts:
  • in correspondence with the startup at time $t_{0}$, we set 5 time steps in an irregular, ascendant order between $t_{0}$ and $t_{0}+2$ d, namely: $t_{0}$, $t_{0}+0.1$, $t_{0}+0.2$, $t_{0}+0.5$, $t_{0}+1$ and $t_{0}+2$. In this way we take into account the buildup of those fission products that have a large absorption cross section (the so called poisons, e.g. the ${}^{135}\text{Xe}$);
  • from $t_{0}+2$ d up to the shutdown time $t_{d}$, we use a uniform time step length of $\Delta t=50$ d.
We run Serpent simulations using an implicit scheme (Stochastic Implicit Euler, SIE) in order to avoid spatial instabilities that occur in 3D burnup calculations BurnupMethods . This method, in a similar way to all implicit schemes, solves the differential equation adopting the (unknown) end of step (EOS) flux value. At the beginning of each burnup step, a predictor passage of Forward Euler is applied. Then, several corrector iterations take place. Each iteration adopts the EOS flux from the previous corrector step. After repeating this procedure a fixed number of times, the final estimation of the nuclide density is obtained with the average of the corrector fluxes. More details about the SIE method can be found in SerpentSIE ; CastagnaImplicit ).
Figure 5: Geometries adopted for the description of fuel rods in the Takahama-3 reactor. (a) 30 burnup zones at the same temperature; (b) 30 axial burnup zones with 5 temperature zones for fuel and 3 temperature zones for water; (c) One burnup zone at uniform temperature. For consistency with respect to the experimental report the z-axis starts from the top of the fuel active region in the downward direction.
In order to map the evolution of nuclide concentrations along the axial direction and compare them with the experimental data, we divided the fuel pins into 30 burnup zones along the z axis. This means that in each volume the fuel burnup is computed starting from the neutron flux evaluated in it. Moreover, in order to estimate the effect of the temperature field characterizing the reactor core, we implemented three different simulations:
  1. (a)
    the same temperature $T_{\text{c,f}}=900$ K is applied to all 30 zones and water is set at $T_{\text{c,w}}=557$ K everywhere.
  2. (b)
    the 30 burnup zones are grouped into 5 macro-zones, applying respectively the temperatures of 750 K, 900 K, 1200 K, 900 K, 750 K. This profile is intended to reflect a realistic temperature distribution in the fuel, consistent with the shape of the neutron flux in Fig. 3, while preserving the average fuel temperature of 900 K. The water between the pins is divided in three zones with increasing temperatures from bottom to top (550 K, 557 K and 563 K).
  3. (c)
    a single burnup zone at uniform temperature is simulated for the whole assembly.
The last case study (c) is not used for the purpose of the experimental benchmark due to the lack of spatial meshing, but is intended to investigate the role of spatial discretization in the computation of fission fractions. Fig. 5 shows a schematic of the three simulated models.
3.5 Results
Figure 6: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF95 pin, compared with simulation results obtained with uniform temperature field (blue) and different temperature zones (red). Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF95 rod.
Figure 7: Top panels: measured axial concentrations of the four fissile isotopes (black dots) in the SF97 pin, compared with simulation results obtained with uniform temperature field (blue) and different temperature zones (red). Bottom panels: C/E ratio of calculated (C) concentrations versus experimentally measured (E) ones for SF97 rod.
The top panels of Figures 6 and 7 show the comparison between the measured axial concentration of the four fissile nuclides against the Serpent predictions. In these plots we also compare the simulation results obtained with the two models for the temperature field: (a) vs (b). Temperature acts directly on reaction rates and, thus, has an impact on fuel burnup. Particularly, at higher temperatures, the neutron flux Maxwellian distribution in the thermal range is peaked at higher energies where fission and capture cross sections are lower, but at the same time the Doppler broadening of cross section resonances has the effect of increasing the rates of reactions induced by neutrons in the intermediate range of the spectrum. Moreover, the temperature and, thus, the density of water has an impact on its moderation capability, which is more effective at lower temperatures. Depending on which of these effects is dominant, the reaction rates will be higher or lower as a function of the material temperature. For example, we observe that ${}^{239}$u accumulates more in the central part of the pin when this is simulated at a higher temperature. This is due to the fact that ${}^{239}$u is produced from neutron captures by ${}^{238}$, which mostly occur at the resonances.
In the bottom panels of Figures 6 and 7 we show the Calculated-to-Experimental (C/E) ratios, i.e. the ratios between the nuclide concentrations predicted by Serpent simulations and those measured experimentally. The closer the C/E ratios are to 1, the more accurate the prediction of nuclide concentration is. In principle, the uncertainty to be associated to C/E ratios is given by the propagation of both experimental and simulation uncertainties. The uncertainties reported in literature associated to the IDMS measurements are $<0.1\%$ for ${}^{235}$ and ${}^{238}$, and $<0.3\%$ for ${}^{239}$u and ${}^{241}$u. Since we run the Monte Carlo simulations with a relatively high number of neutrons (150 active cycles of 5×$\times 10^{10}$ neutrons/cycle), the statistical uncertainty of MC results is negligible with respect to that associated to the experimental data.
Looking at the C/E results shown in Figures 6 and 7, a very good agreement within ±$\pm 0.3\%$ is found, as expected, for ${}^{238}$, whose concentration remains very similar to the fresh fuel value. The other three fissile isotopes (${}^{235}$, ${}^{239}$u, and ${}^{241}$u) exhibit differences within the range $\pm$10in all sampling sites except for the uppermost ones (SF95-1, SF97-1, and SF97-2). By comparing the simulations performed with the two models of the temperature field (uniform vs axial profile), we observe local effects increasing or decreasing the nuclide concentrations up to 0.7%. Thus, the temperature field has a relatively small impact if compared with the systematic differences between experimental data and simulations.
The data in the top locations show a worse agreement with the simulation results probably because these samplings were made close to the upper end of the fuel pin. In fact, the estimated uncertainty in the sampling position is $<2$ cm Suyama2008 . This is a negligible error in the central pin zone, where the flux profile becomes flat as the depletion goes on. On the contrary, at the pin edges where the neutron flux has a significant gradient, a 2 cm shift can introduce a non-negligible systematic error. Moreover, the neutronics at the axial edges depends on the back-scattering capability of the reflector. Unfortunately, the Takahama-3 reactor design data for reflector and cladding cap length and composition are not detailed in the technical reports. Their dimensions and compositions were hypothesised according to Suyama2008 , thus we simulated 2 cm of zircaloy-4 for cladding caps and a 20 cm layer of water.
For this reason, the following data analysis does not take into account the sampling performed in SF95-1, SF97-1 and SF97-2. Table 4 reports the average C/E for the considered samples (SF95 2-5 and SF97 3-6). In this table we also report, for comparison, the results published in literature obtained with other simulation tools (evaluating the average C/E avoiding the values at the rod extremes), namely: SWAT and ORIGEN 2 PIE_ORNL , SCALE 5.1 SCALE_PIE_ORNL and SCALE 4.4 and HELIOS Takahama3 . A similar analysis was previously conducted by Takahama_DRAGON_MURE , comparing a deterministic code (DRAGON) and a Monte Carlo (MURE).
The results obtained in this work with Serpent simulations are characterized by similar C/E values when compared with those obtained in other analyses. For what concerns the excluded data (i.e. SF95-1, and SF97 1-2), results from literature agree with our analysis, with differences with respect to unity up to 35%.
Rod Code ${}^{235}$ ${}^{238}$ ${}^{239}$u ${}^{241}$u
SF95 Serpent (a) 1.061 0.994 1.029 0.979
Serpent (b) 1.0422 1.0002 1.032 0.9577
SWAT 1.015 1.0 1.009 0.959
ORIGEN 2.1 1.016 1.0 1.008 0.976
HELIOS 1.008 1.0 1.026 1.01
SCALE 4.4 0.985 1.0 0.976 0.95
SCALE 5.1 1.028 1.002 1.097 1.017
SF97 Serpent (a) 1.024 0.999 0.984 0.995
Serpent (b) 0.995 0.999 0.958 0.972
SWAT 1.015 1.0 1.031 0.986
ORIGEN 2.1 1.032 1.0 1.050 1.019
HELIOS 1.01 1.0 1.03 1.025
SCALE 4.4 0.97 1.0 0.993 0.975
SCALE 5.1 1.02 0.999 1.048 0.975
Table 4: Average Serpent C/E for chosen samples (SF95 2-5 and SF97 3-6) in the uniform temperature case (a) and with the model with temperature zones (b), compared with the literature results obtained with different simulation codes.
4 Fission fractions and $\bar{\nu}_{e}$ spectra results
Figure 8: Top: Fission fraction evolution in the SF95 rod predicted by simulation (a) with uniform temperature field and simulation (b) with non-uniform temperature field. Bottom: Relative fission fraction differences between simulation (b) and simulation (a) with respect to (b).
Figure 9: Top: Fission fraction evolution in the SF97 rod predicted by simulation (a) with uniform temperature field and simulation (b) with non-uniform temperature field. Bottom: Relative fission fraction differences between simulation (b) and simulation (a) with respect to (b).
In the framework of reactor $\bar{\nu}_{e}$ spectrum prediction, the estimation of the fission fractions of fissile nuclides is more relevant than their concentrations. In this section, first we present the time evolution of fission fractions obtained with different levels of approximation (Sect. 4.1). Then, we show the non-oscillated $\bar{\nu}_{e}$ spectrum, at different fission fractions values (Sect. 4.2).
4.1 Fission Fractions
In the top pads of Figures 8 and 9 we show the time evolution of the fission fractions in the Takahama-3 SF95 and SF97 rods, computed with Serpent simulations using uniform (a) or non-uniform (b) temperature field, as described in Sect. 3.4. With fresh fuel, the contribution to the fission events comes from ${}^{235}$ ($\sim$ 94%) and ${}^{238}$ ($\sim$ 6%). As the fuel depletion proceeds, two major effects can be appreciated: ${}^{235}$ is consumed and Pu nuclides are produced from ${}^{238}$ neutron captures. Therefore, with the increase in the burnup level, the relative importance to the total fission events of ${}^{235}$ tends to decrease, while the ${}^{239}$u and ${}^{241}$u contribution increases. The grey areas represent the cooling periods between cycles in which only radioactive decay occurs. In the bottom pads of Figures 8 and 9 we show the relative differences of fission fractions between simulation (a) and simulation (b). For both SF95 and SF97 rods, the relative differences have similar trends, with a weak dependence on burnup for ${}^{235}$, ${}^{238}$, ${}^{239}$u. For ${}^{241}$u we observe an unstable trend at the beginning of fuel depletion, which can be considered not relevant in our analysis since its fission fraction value is nearly zero at low burnup. It is worth noting here that on average, after the first operational cycle, the fission fractions predicted by simulation (b) differ from those predicted by simulation (a) by about:
  • (${}^{235}$) $+0.85\%$ in SF95 and $+0.23\%$ in SF97 rod;
  • (${}^{238}$) $-9.0\%$ in both SF95 and SF97 rods;
  • (${}^{239}$u) $+0.34\%$ in SF95 and $+0.92\%$ in SF97 rod;
  • (${}^{241}$u) $+1.1\%$ in SF95 and $+2.4\%$ in SF97 rod;
In light of these findings, we can state that modelling the temperature field in MC simulations has a non-negligible effect in the estimation of fission fractions, especially the ${}^{238}$ one. The temperature itself has a lower impact on the fissions from the ${}^{238}$. The relative reduction in the ${}^{238}$ fraction is addressed to the broadening of the fission cross sections resonances of the other nuclides (${}^{235}$, ${}^{239}$u, and ${}^{241}$u) in the hotter fuel volumes where the neutron flux is more intense. Thus, since the sum of fission fractions is normalized to 1, increasing the fission probability of the other three nuclides causes a decrease of the ${}^{238}$ fraction. Therefore, this analysis highlights that an overly approximate simulation model that does not take into account the temperature field can lead to fission fraction estimates affected by systematic errors on the scale of the differences listed above.
Finally, we also analyzed the fission fraction differences between the simulation (a), with respect to the one with a single burnup zone (simulation (c)). In this case the differences are $\lesssim 2\%$ for all nuclides.
4.2 Non oscillated $\bar{\nu}_{e}$ spectra results
Figure 10: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the fission fractions corresponding to three different burnup levels of SF97 rod. Bottom: Relative differences of spectra after 550 d and 1340 d with respect to the fresh fuel spectrum.
Figure 11: Top: $\bar{\nu}_{e}$ spectra multiplied by the IBD cross section computed with the fission fractions at the end of the burnup cycle for the three simulations of the SF95 rod. Bottom: Relative differences of spectra from simulations (b) and (c) with respect to simulation (a).
We finally present the equilibrium $\bar{\nu}_{e}$ spectra obtained through the summation method by combining the cumulative fission yield (CFY) data published in the Live Chart of Nuclides by IAEA IAEA_livechart and the $\bar{\nu}_{e}$ spectra of 468 $\beta^{-}$ decays obtained through BetaShape by setting an energy binning of 1 keV. We use thermal CFY for ${}^{235}$, ${}^{239}$u and ${}^{241}$u, which are mostly fissioned by thermal neutrons, and fast CFY for ${}^{238}$ which can be fissioned with not negligible probability by fast neutrons only. Then, the total reactor $\bar{\nu}_{e}$ spectrum (Equation (2)) is evaluated with the fission fractions estimated with Serpent simulations. Since the reactor $\bar{\nu}_{e}$ will be detected through the IBD reaction, we multiply the $\bar{\nu}_{e}$ spectra by the IBD cross section STRUMIA200342 so as to obtain the shape of the detectable non oscillated spectrum. We want to highlight that this analysis is meant to compare the propagation of the reactor model differences on the reactor antineutrino spectrum. Other reactor antineutrino calculation methods as well as comparison with experimental data is out of the scope of the analysis and will be presented in future works.
In Fig. 10, we show the $\bar{\nu}_{e}$ spectra obtained by using the fission fractions estimated by Serpent for the Takahama-3 assembly at 3 different burnup levels for the SF97 rod: fresh fuel, intermediate burnup fuel (after 550 days), and end-of-life fuel (1340 days). As shown in the bottom pad, the $\bar{\nu}_{e}$ flux intensity decreases with time. This is because the average number of $\bar{\nu}_{e}$ emitted after ${}^{239}$u fission is lower with respect to that of ${}^{235}$. The relative difference calculated with respect to the fresh fuel $\bar{\nu}_{e}$ flux is of the order of 10% around the maximum of the spectrum at 4 MeV, and is up to $\sim 30\%$ in the high energy range. This result reflects the extreme case where the reactor starts from the fresh condition, without any refuelling strategy: in a more realistic scenario, this effects is averaged out considering fuel at different depletion stages. Nevertheless, in the framework of the JUNO experiment, it will be very important to correctly sum up the non-oscillated spectra taking into account the fission fraction evolution of each of the eight reactors during the data taking.
Finally, in Fig. 11 we show the impact of the systematic errors discussed in Sect. 4.1 on the non oscillated $\bar{\nu}_{e}$ spectrum. In particular, the spectra shown in the top pad were computed by using the fission fractions obtained from the simulations (a), (b) and (c) of the SF95 rod at the end of its second cycle (875 days). In this case, the differences are much less pronounced, remaining below 1% up to 7 MeV and in any case below 3% at higher energies. This can be addressed to the fact that the major differences are related to the ${}^{238}$ fraction, which has a minor impact on the overall system.
5 Conclusions
In this paper we presented a methodology for the prediction of the non-oscillated reactor $\bar{\nu}_{e}$ spectrum. This is intended to serve the scientific community in future reactor $\bar{\nu}_{e}$ experimental campaigns (e.g., JUNO and TAO) that will analyze the mass ordering problem with an unprecedented level of precision. We coupled the Monte Carlo code Serpent along with the $\bar{\nu}_{e}$ spectrum calculation software BetaShape. The Serpent simulations allow the calculation of fission fraction evolution over time, leveraging data encompassing reactor geometry, power history, and fresh fuel inventory. Subsequently, by integrating BetaShape alongside nuclear databases, we construct the antineutrino spectra emanating from the four primary fissile isotopes with the summation method – ${}^{235}$, ${}^{238}$, ${}^{239}$u, ${}^{241}$u – thereby enabling predictions of the total reactor antineutrino spectra across various burnup levels.
In this paper we focused our analysis on the fuel burnup, aiming at exploring the accuracy in the evaluation of fission fractions through MC simulations. For this purpose, we used the spent fuel concentrations of the Takahama-3 PWR as experimental benchmark for the Serpent simulations. We found that Serpent simulation run with either uniform temperatures in the fuel/moderator or imposing different temperature zones, are both in a reasonably good agreement with the experimental data. Importantly, our simulations highlight that the temperature field introduces minor discrepancies in the fuel depletion calculations, its effect but its influence is more pronounced when considering the fission fractions. Looking ahead, our future plans encompass harnessing this analytical tool to further investigate additional sources of systematic uncertainty (such as cross-section or fission yield data uncertainties) and their subsequent impact on the antineutrino spectrum. Future developments will be devoted to relax the equilibrium hypothesis on the fission products, highlighting the importance that this amount of nuclide have on the spectrum. Also, a more detailed analysis of the reactor configuration of the Taishan and Yangjiang plant will be addressed, along with the comparison with other available experimental spectra from near-detector experiments.